# EquiAV: Leveraging Equivariance for Audio-Visual Contrastive Learning

### v1

@@ 4.3 Qualitative Results
The evaluation results and ablation studies indicate that the proposed representative embedding effectively captures the salient information of audio-visual inputs. We also explore how the relationship between the original embedding and the representative embedding is manifested in the input space. To this end, 1,000 augmentation vectors are randomly generated for each sample and then the equivariant embedding nearest to the centroid is selected as the substitute for the centroid. The corresponding transformation for this equivariant embedding is likely the one that best represents the image in the input space. The qualitative results in Figure 3 further validate our hypothesis, specifically by highlighting the parts of the image corresponding to the paired audio. In addition, the zero-shot sound source localization task is performed on the original images and the images selected by the above method. While the attention map on the original images is dispersed to parts beyond the main elements, the attention for the augmented images, which correspond to the substitute embeddings of the centroids, is precisely and intensively concentrated on the areas that make the sound.
@@ 5 Conclusion
Appendix A Analysis on Equivariant Loss Functions
As mentioned in ablation study on Table 8, our equivariant loss function differs from the equivariant contrastive loss function used in EquiMod (Devillers & Lefort, 2023), regarding whether a positive pair similarity is included in the denominator. The equivariant loss of EquiMod is the same as applying an indicator function to both summation terms in the denominators of Equation 9, which can be represented as follows:
$\ell_{\text{EquiMod}}(\hat{z},z^{\prime})=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\hat{z}_{i},{z^{\prime}}_{i})/\tau)}{\sum_{j=1}^{N}\mathbbm{1}_{[j\neq i]}\left[\exp(\text{sim}(\hat{z}_{i},{z^{\prime}}_{j})/\tau)+\exp(\text{sim}(\hat{z}_{i},\hat{z}_{j})/\tau)\right]}$ (A)
@@ Appendix A Analysis on Equivariant Loss Functions
Appendix B Implementation Details
@@
  • Gaussian blur (1 element for both audio & visual): the value of $\sigma$ for Gaussian blurring kernel (0 as default).
@@
Appendix C Hyperparameter Details
The hyperparameter settings used in this paper are listed in Table A.
Table A: Hyperparameters used in pre-training and fine-tuning phase.
Stage Pre-training Fine-Tuning
Dataset AudioSet-2M AudioSet-20K AudioSet-2M VGGSound ESC-50 SPC-v1
Optimizer AdamW
Optimizer momentum $\beta_{1}$=0.9,  $\beta_{2}$=0.95
Weight decay 1e-5 1e-5 1e-5 1e-5 1e-5 1e-5
Learning rate scheduler half-cycle cosine annealing (Loshchilov & Hutter, 2016)
Initial learning rate 1e-6 1e-6 1e-6 1e-6 1e-6 1e-6
Peak learning rate 1e-4 1e-4 1e-4 1e-4 5e-4 1e-3
Warm-up epochs 2 1 1 1 4 1
Epochs 20 50 50 50 60 10
Batch size 256 256 512 256 64 256
Class Balancing Weight No No Yes Yes No No
Mixup No Yes Yes Yes No Yes
Loss Function EquiAV Loss (Eq. 16) BCE BCE CE CE BCE
Temperature ($\tau$) 0.07 - - - - -
Input Norm Mean -4.346 -4.346 -4.346 -4.956 -6.627 -6.702
Input Norm STD 4.332 4.332 4.332 4.486 5.359 5.448
GPUs 8 A6000 8 A5000 8 A6000 8 A5000 8 A5000 8 A5000

### v2

@@ 4.3 Qualitative Results
The evaluation results and ablation studies indicate that the proposed representative embedding effectively captures the salient information of audio-visual inputs. We also explore how the relationship between the original embedding and the representative embedding is manifested in the input space. To this end, 1,000 augmentation vectors are randomly generated for each sample and then the equivariant embedding nearest to the centroid is selected as the substitute for the centroid. The corresponding transformation for this equivariant embedding is likely the one that best represents the image in the input space. The qualitative results in Figure LABEL:fig:qual further validate our hypothesis, specifically by highlighting the parts of the image corresponding to the paired audio. In addition, the zero-shot sound source localization task is performed on the original images and the images selected by the above method. While the attention map on the original images is dispersed to parts beyond the main elements, the attention for the augmented images, which correspond to the substitute embeddings of the centroids, is precisely and intensively concentrated on the areas that make the sound.
@@ 5 Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgements
This work was supported in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT, No. 2022-0-00184, Development and Study of AI Technologies to Inexpensively Conform to Evolving Policy on Ethics), by the National Research Foundation of Korea grant funded by the Korean government (MSIT, RS-2023-00212845), and by the Information Technology Research Center (ITRC) support program (IITP-2024-RS-2023-00259991) supervised by Institute of Information & communications Technology Planning & Evaluation (IITP).
Appendix A Appendix
Appendix B Additional Ablation Studies
Weight-Scales of Loss Function.
To understand how audio-visual contrastive learning is influenced by learning from each data modality, we explored the effects of setting $\lambda_{a,intra}$ and $\lambda_{v,intra}$ in Table A. When self-supervision is derived solely from the audio domain (i.e. $\lambda_{v}=0$), it yielded the highest performance in audio-only tasks but showed suboptimal performance in visual-only and audio-visual tasks, and vice versa. Furthermore, adjusting the values of $\lambda_{a}$ and $\lambda_{v}$ to be either higher or lower relative to $\lambda_{inter}$ also resulted in relatively lower performance. Consequently, we found that using equal $\lambda$ values for both modalities and inter-modality learning produced balanced results.
Table A: Audio-visual event classification performance on AudioSet-20K with varying weights scales of loss function.
Fine-Tuning $\lambda_{inter}$ $\lambda_{a;intra}$ $\lambda_{v;intra}$ A V A-V 1 1 0 42.9 24.0 43.4 1 0 1 41.2 26.3 43.2 1 2 2 42.2 25.5 45.5 1 0.5 0.5 41.3 25.1 46.0 1 1 1 42.4 25.7 46.6
Computational Cost & Scalability.
Table B reports the computational costs required by generating equivariant representations and Table C shows how the size of the pre-training dataset affects performance on downstream tasks. The results demonstrate the computational efficiency and scalability of our framework.
Table B: Computation cost for generating equivariant representations in audio and visual domains.
##\## representations Audio(GFLOPs) Visual(GFLOPs) Baseline 97.80 35.20 1 98.21 (+0.41) 35.36 (+0.16) 4 98.23 (+0.43) 35.38 (+0.18) 8 98.25 (+0.45) 35.40 (+0.20) 16 98.30 (+0.50) 35.44 (+0.24)
Table C: Zero-shot retrieval results on AudioSet and audio-visual event classification performance on AudioSet-20K with the variants of pre-training datasets.
Zero-shot Retrieval Fine-Tuning (mAP) Pre-training Dataset V2A A2V A V A-V AudioSet-20K 3.9 4.2 33.7 17.3 35.2 VGGSound 11.2 10.5 37.6 21.3 40.8 AudioSet-2M 30.1 29.6 42.4 25.7 46.6
Appendix C Analysis on Equivariant Loss Functions
As mentioned in the ablation study on Table 9, our equivariant loss function differs from the equivariant contrastive loss function used in EquiMod (Devillers & Lefort, 2023), regarding whether a positive pair similarity is included in the denominator. The equivariant loss of EquiMod is the same as applying an indicator function to both summation terms in the denominators of Equation 9, which can be represented as follows:
$\ell_{\text{EquiMod}}(\hat{z},z^{\prime})=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(\hat{z}_{i},{z^{\prime}}_{i})/\tau)}{\sum_{j=1}^{N}\mathbbm{1}_{[j\neq i]}\left[\exp(\text{sim}(\hat{z}_{i},{z^{\prime}}_{j})/\tau)+\exp(\text{sim}(\hat{z}_{i},\hat{z}_{j})/\tau)\right]},$ (A)
@@ Appendix C Analysis on Equivariant Loss Functions
Appendix D Implementation Details
@@
[…]
@@
Appendix E Hyperparameter Details
The hyperparameter settings used in this paper are listed in Table D.
Table D: Hyperparameters used in pre-training and fine-tuning phase.
Stage Pre-training Fine-Tuning Dataset AudioSet-2M AudioSet-20K AudioSet-2M VGGSound ESC-50 SPC-v1 FSD-50K UCF101 HMDB51 K400 Optimizer AdamW Optimizer momentum $\beta_{1}$=0.9,  $\beta_{2}$=0.95 Weight decay 1e-5 Learning rate scheduler half-cycle cosine annealing (Loshchilov & Hutter, 2017) Initial learning rate 1e-6 Peak learning rate 1e-4 1e-4 1e-4 1e-4 5e-4 1e-3 5e-4 1e-3 5e-4 5e-4 Warm-up epochs 2 1 1 1 4 1 4 0 0 4 Epochs 20 50 50 50 60 10 60 100 100 100 Batch size 256 256 512 256 64 256 64 32 32 512 Class Balancing Weight No No Yes Yes No No No No No No Mixup No Yes Yes Yes No Yes No No No Yes Loss Function EquiAV Loss (Eq. 16) BCE BCE CE CE BCE BCE CE CE CE Temperature ($\tau$) 0.07 - - - - - - - - - Input Norm Mean -4.346 -4.346 -4.346 -4.956 -6.627 -6.702 -6.627 - - - Input Norm STD 4.332 4.332 4.332 4.486 5.359 5.448 5.359 - - - GPUs 8 A6000 8 A5000 8 A6000 8 A5000 8 A5000 8 A5000 8 A5000 8 A5000 8 A5000 8 A5000